OpenStudy (anonymous):

5x - 3y = 15 oredered pair by ( ,5),( ,0),(2, ) x + 4y = -2 oredered pair by (0, ),( ,0),(1 , )?

OpenStudy (anonymous):

i got you r back ok v

OpenStudy (anonymous):

3x + 4y = 27 5x - 3y = 16 3x + 4y = 27 5(3x + 4y) = 5(27) 15x + 20y = 135 5x - 3y = 16 3(5x - 3y) = 3(16) 15x - 9y = 48 ....15x + 20y = 135 –) 15x ...- 9y = ..48 ----------------------------------- .............29y = ..87 29y = 87 y = 87/29 y = 3 3x + 4y = 27 3x + 4(3) = 27 3x + 12 = 27 3x = 27 - 12 x = 15/3 x = 5 ∴ x = 5 , y = 3

OpenStudy (triciaal):

when x = 0 solve for y 5(0)-3y = 15 -3y = 15 divide each side by -3 y = -5 when y = 0 5 x -3(0) = 15 so x = 5

OpenStudy (anonymous):

help me to answer my question plz.